part1. The Psych 230 students finished the course with a total population mean of 155 points...

part1.

The Psych 230 students finished the course with a total population mean of 155 points and the standard deviation is 50 points. One student earned 300 points. What is the z-score for this raw score?

z-score =

part 2.
The average feet for American men are normally distributed, with a mean of 11 inches and a standard deviation of 0.66 inches. The average feet for American women are also normally distributed, but with a mean of 10 inches and a standard deviation of 0.56 inches.

a) If a man's feet are 11 inches long, what is his z-score (to two decimal places)?
z =
b) If a woman's feet are 9.7 inches long, what is her z-score (to two decimal places)?
z =
c) Which is relatively longer?

  • The woman's feet
  • The man's feet

part 3
Find the probability distribution of the variable for number of assessments a student from the class has failed in.

x P(x)
0 0.05
1 0.3
2 0.2
3 0.45



Find the mean of this probability distribution. Round your answer to two decimal places.
Answer #1

Part 1:

Z = (300 - 155)/50 = 2.9

So,

Z- score = 2.9

Part 2:

(a)

Men:

Z = (11 - 11)/0.66 = 0

So,

Z- score = 0

(b)

Women:

Z = (9.7 - 10)/0.56 = - 0.5357

So,

Z- score = - 0.5357

(c)

Correct option:

The man's feet

Part 3:

x p xp
0 0.05 0
1 0.3 0.3
2 0.2 0.4
3 0.45 135
Total = 2.05

So,

mean of this probability distribution = 2.05